10 men can dig a piece of land in 12days. How many more men are required to dig it for 4days?
Answer by ankor@dixie-net.com(22740)

10 men can dig a piece of land in 12days
Find how many man-days required to do the job
10 * 12 = 120 man-days
:
How many more men are required to dig it for 4days?
let m = no. of men required to do the same job in 4 days
4m = 120
m = 120/4
m = 30 men total required
But they ask how many MORE
30 - 10 = 20 more men